# The SHMS 11 GeV/c Spectrometer in Hall C at Jefferson Lab

### v1

@@ 4.1.1 Design and Construction
To meet the requirements listed above, the SHMS Hodoscope was built as a series of arrays (planes) of plastic scintillator paddles. The S1X and S1Y planes have 13 1008 cm paddles each, while the S2X plane has 14 11010 cm paddles. For each of the three scintillator planes, the paddles were staggered in the beam direction by 0.7 cm and overlapped transversely by 0.5 cm. To minimize the impact of the scintillators on downstream detectors and also to ensure good timing resolution, the thickness of paddles was 0.5 cm.
The scintillator material used was Rexon RP-408 [41]. The paddles were wrapped by the manufacturer with millipore paper, aluminum foil, and 2” wide electrical tape. The transition between the thin scintillator material and the photomultiplier (PMT) tubes was done using a Lucite™ fishtail-shaped light guide. As the glued joint between the scintillator paddle and the light guide is rather fragile (0.58.0 and 0.510.0 cm joints, respectively) aluminum “splints” were used to reinforce it. The PMT to fishtail joint was originally wrapped with 2” tape as well and light-leak tested; subsequently this wrapping was reinforced with TEFLON™ tape and a 3” heat-shrink sleeve.
Each scintillator paddle has a PMT at each end glued to the fishtail using optical glue (BC-600 [42]) matching the index of refraction of the Lucite™. A combination of Photonis XP 2262[43] and ET 9214B [44] 2” tubes were used. Both models have 12-stage amplification and their maximum photocathode sensitivity is in the blue–green range. The typical gain is . Gains were measured as a function of high voltage during the construction and the whole hodoscope was gain matched in situ once installed in SHMS.
@@ 4.2.1 Design and Construction
Quartz bars of 2.55.5125 cm3 dimensions with an index of refraction of approximately 1.5 were procured from Advanced Glass Industries. The Cherenkov light produced by electrically charged particles is detected by UV-glass window PMTs (model ET9814WB) and quartz window ET9814QB [45] photomultiplier tubes optically coupled to the quartz bars through RTV™615 silicon rubber of 50 m thickness. Thicker optical couplings resulted in fewer photoelectrons, so excellent quality control of the maximum thickness was required. The 16 bars in use in the hodoscope quartz plane are staggered in the beam direction with a transverse overlap between adjacent bars of 0.5 cm. The quartz plane frame allows for more bars to be added.
4.2.2 Calibration
@@ 4.3 Drift Chambers
4.3.1 Design
The SHMS horizontal drift chambers provide information to determine the trajectory of charged particles passing through the detector stack. The drift chamber package consists of two horizontal drift chambers separated by a distance of 1.1 m and oriented in the detector stack such that the sense wires planes are perpendicular to the central ray. Each chamber consists of a stack of six wire planes providing information on the track position along a single dimension in the plane of the wires and perpendicular to the wire orientations to better than 250 m. The perpendicular distance of the track relative to the wire is determined from the time of the signal produced by the ionization electrons as they drift from their production point to the wire in an electric field of approximately 3700 V/cm.
@@ 4.3.1 Design
Each chamber consists of two identical half chambers separated by a fiberglass mid-plane which supports the amplifier discriminator cards required for the sense wire readout. To minimize the production costs, only two unique PCB types were designed: an X-plane with wires oriented horizontally (left panel of Figure 26), and a U-plane with wires oriented at +60∘ relative the X-plane (middle panel of Figure 26). All other plane orientations are generated by rotations of these two basic board types. For instance, the boards are designed such that a rotation of 180∘ in-plane about an axis through the center of the board produces boards with wires of the same orientation, but shifted by 1/2 cell width, thus allowing the resolution of left/right ambiguities. Rotation of Figure 26 such that the top becomes the bottom produces the X’ and U’ orientations. The V and V’ boards, with wire orientation of -60∘ relative to the X-plane, are produced by a rotation of the U and U’ boards of 180∘ into the page about a vertical axis though the center of the board. Each half chamber has three planes with the first half consisting of (U, U’, X) and the second half consisting of (X’, V’, V). The first chamber is oriented in the SHMS frame such that the board ordering as seen by particle traversing the spectrometer is (U, U’, X, X’, V’, V), while for the second chamber the ordering is reversed (V, V’, X’, X, U’, U). A drawing showing the chambers mounted in the frame is presented in Figure 27.
@@ 4.3.2 Calibration
In the approximation that the incident particle distribution is uniform, or has a constant slope, a single cell222A cell is one sense wire surrounded by field wires such that the sense wire is at the center and the field wires are at the corners. on average sees a flat distribution of events. The measured drift time distribution can therefore be mapped to drift distance by areal scaling. This distribution can be determined for the individual wire or averaged over an entire group (up to 16 wires per discriminator card). Associated with each drift time distribution is a time, , which corresponds to the time at which ionized particles come into contact with the wire. If this value is non-zero, this is the value by which all drift times must be shifted in order to assure that . All subsequent times in each spectrum are measured relative to this time.
@@ 4.4 Heavy-Gas Cherenkov Counter
4.4.1 Design
The SHMS Heavy-Gas Cherenkov detector (HGC) is a threshold-type Cherenkov detector, designed to separate charged and over most of the SHMS operating momentum range, 3–11 GeV/c. The radiator gas C4F10 at 1 atm, with an index of refraction of =1.00143 at standard temperature [46], allows to produce abundant Cherenkov light above 3 GeV/c momentum, while remain below Cherenkov threshold until about 7 GeV/c. Optimal separation at higher momenta requires a reduction in the gas pressure, down to 0.3 atm at 11 GeV/c.
A schematic view of the detector is shown in Fig. 29. The SHMS focal plane is subtended by four 5560 cm 0.3 cm thick glass mirrors, which reflect the Cherenkov radiation to four Hamamatsu R1584 12.5 cm diameter photomultiplier tubes located above and below the particle envelope. The mirrors and gas are enclosed in a cylindrical aluminum tank of 164.9 cm inner diameter and 113.5 cm length, with entrance and exit windows of hydroformed 0.102 cm thickness 2024 T-4 aluminum alloy [47]. The vessel is sufficiently strong to be pumped to vacuum before introducing the radiator gas, avoiding the need to purge when filling. A unique aspect of the detector is the placement of the photomultipliers outside the gas envelope, viewing the enclosure through 1.00 cm thick Corning 7980 quartz windows. This allows the gas enclosure to be smaller in diameter than would otherwise be possible, as the full length of the PMT and base no longer need to be fully within the diameter of the vessel. It also makes the PMTs available for servicing without venting the gas.
The mirrors are inexpensive, having been produced by the slumping process [48]. As a result, they deviate from the desired 110 cm radius of curvature with a slightly oblate shape [49]. However, the Cherenkov cone on the mirrors for 3-7 GeV/c in C4F10 is 7-10 cm in diameter, so optical quality mirrors are not required for this application. The UV wavelength characteristics of the respective optical components are relatively well matched. C4F10 has good transmittance down to 160 nm [46]. The quartz viewing windows provide 88% transmission down to 200 nm, including the 10% loss due to surface reflection [50], and the optical glass faced PMTs have 70% of their peak quantum efficiency at 200 nm (peak at 350 nm) [51]. Accordingly, the mirror reflectivity was optimized for 90% at 270 nm, and 75% at 200 nm [52].
@@ 4.4.2 Calibration
where is the distance from the focal plane to the HGC mirror plane. The coordinate axis for the HGC is the convention used in charged particle transport in dispersive magnetic systems. The -axis is the direction of increasing particle momentum, the -axis is the direction of particle travel through the spectrometer, and the -axis is deduced from . Additionally, timing cuts are applied to the HGC data, collected using the high resolution pulse time setting in the FADC250’s FPGA [53]. The pulse time reconstruction in firmware corresponds to the time at which a pulse reaches half of its maximum amplitude after passing a threshold of 5 mV. Lastly, a cut on particle velocity, , is applied.
An example of a completed calibration is shown in Figs. 30, 31. For this run, the HGC was filled with C4F10 at 1 atm, and the SHMS central momentum was 2.583 GeV/c with polarity set to detect positively-charged particles. Cherenkov radiation is produced by traversing the HGC with momentum GeV/c. This can occur only for , which corresponds roughly to the bottom half of the HGC. Subthreshold with , as well as and , may produce low-level light in the HGC via knock-on electron emission and scintillation in the radiator gas. The adjacent mirror cuts described above produce a clear SPE peak in Fig. 30.
@@
An inherent systematic uncertainty is present in the HGC calibration due to statistical errors in determining the location of the SPE peak in the various mirror quadrants. This uncertainty was quantified by recording the locations of the SPE across several runs, for the different adjacent mirror combinations for each PMT, as well as by varying the contribution of the higher PE tail extending underneath the SPE peak, as in Figs. 30, 31. The systematic uncertainty in the calibration is taken to be the root mean square of this set of values, giving %. It should be noted this uncertainty is somewhat larger than the statistical uncertainty of the SPE peak, which is typically 0.2 to 0.6%.
@@ 4.5.1 Design
Four spherical thin glass mirrors of radius 135 cm, square in shape with edge lengths of 43 cm, focus the Cherenkov light onto the PMTs. The glass blanks were manufactured by Rayotek Scientific [54] from borosilicate glass of 3 mm thickness by slumping over a polished steel mold and then cutting to dimensions. Simulations showed a reduction of collection efficiency due to incoming photon losses at the exposed edges of the mirror. As such, the edges were bevelled away from the active surface to minimize scattering from these edges.
@@
The blanks were coated by the Thin Film and Glass Service of the Detector Technologies Group at CERN [55]. The reflectivity was also measured at CERN and found to be excellent well into the UV (Fig. 34).
Like the HGC, the four mirrors are arranged in a 2 by 2 array with a small overlap in the center, providing full coverage over the active area. In order to accomplish this without mechanical interference, the mirrors were staggered at slightly different along the tank -axis. The mirrors were mounted in a monolithic frame installed as single unit (see Figure 35).
The four PMTs are 14 stage 5” quartz window PMTs manufactured by Electron Tubes Enterprises [56], model 9823QKB04. The 9823QKB04 has a quantum efficiency above 5% at 150 nm and 30% at 350 nm. The tubes are surrounded by a mu-metal shield and the positive HV is distributed to the stages by a resistive divider. After evidence of corona or arcing inside one of the bases due to the relatively low dielectric breakdown strength of the gas mixture, the problem was resolved by applying silicon conformal coating to suspect areas.
@@ 4.6 Aerogel Cherenkov Counter
4.6.1 General Design Overview
Fig. 37 shows a photograph of the aerogel counter installed downstream of the cylindrical HGC in the SHMS detector stack. The detector consists of two main components: a tray which holds the aerogel material, and a light diffusion box with photomultiplier tubes (PMTs) for light readout. Four identical trays for aerogel of nominal refractive indices of 1.030, 1.020, 1.015 and 1.011 were constructed. These allow for particle identification over the wide range of momenta summarized in Table 4. The design allows for easy detector assembly and replacement of the aerogel trays. Using up to 9 cm aerogel thickness in the trays, the total thickness of the detector is 24.5 cm along the optical axis of the SHMS. A detailed discussion of the detector, characterization of its components, and performance tests can be found in Refs. [57].
@@ 4.6.1 General Design Overview
The trays contain two different types of aerogel. The aerogel tiles of refractive indices =1.030 and =1.020 were manufactured by Matsushita Electric Works, Ltd prior to 2010. The tiles of refractive indices =1.015 and =1.011 were manufactured by the Japan Fine Ceramics Center between 2010 and 2013. All tiles have dimensions of approximately 11 cm by 11 cm by 1 cm and are hydrophobic. The depth of the aerogel radiator in their trays is on average 9 cm thick (8 layers). For the SP-30, SP-20 and SP-15 trays the aerogel covers and area of 110 cm x 100 cm. In the SP-11 aerogel tray the radiator covers an active area of 90 cm x 60 cm. To improve the reflectivity, and thus light collection, inside the trays for the lowest two refractive indices 1 mm thick GORE diffusive reflector material (DRP-1.0-12x30-PSA) with reflectivity of about 99% was used. The SP-30 and SP-20 aerogel trays were covered with 0.45 m thick Millipore paper membrane GSWP-0010 (Millipore).
@@ 4.7.1 Preface
A nearly complete description of the design, pre-assembly component checkout, and construction of the SHMS calorimeter can be found in Ref [58]. This section provides a brief summary and unpublished details.
4.7.2 Construction

### v2

@@ 4.1.1 Design and Construction
To meet the requirements listed above, the SHMS Hodoscope was built as a series of arrays (planes) of plastic scintillator paddles. The S1X and S1Y planes have 13 1008 cm paddles each, while the S2X plane has 14 11010 cm paddles. A sketch of the S1X and S2Y planes is shown in Fig. 24. For each of the three scintillator planes, the paddles were staggered in the beam direction by 0.7 cm and overlapped transversely by 0.5 cm. To minimize the impact of the scintillators on downstream detectors and also to ensure good timing resolution, the thickness of paddles was 0.5 cm.
The scintillator material used was Rexon RP-408 [Rexon_RP-408]. The paddles were wrapped by the manufacturer with millipore paper, aluminum foil, and 2” wide electrical tape. The transition between the thin scintillator material and the photomultiplier (PMT) tubes was done using a Lucite™ fishtail-shaped light guide. As the glued joint between the scintillator paddle and the light guide is rather fragile (0.58.0 and 0.510.0 cm joints, respectively) aluminum “splints” were used to reinforce it. The PMT to fishtail joint was originally wrapped with 2” tape as well and light-leak tested; subsequently this wrapping was reinforced with TEFLON™ tape and a 3” heat-shrink sleeve.
Each scintillator paddle has a PMT at each end glued to the fishtail using optical glue (BC-600 [BC-600]) matching the index of refraction of the Lucite™. A combination of Photonis XP 2262[Photonis_XP2262] and ET 9214B [ET_9214B] 2” tubes were used. Both models have 12-stage amplification and their maximum photocathode sensitivity is in the blue–green range. The typical gain is . Gains were measured as a function of high voltage during the construction and the whole hodoscope was gain matched in situ once installed in SHMS.
@@ 4.2.1 Design and Construction
Quartz bars of 2.55.5125 cm3 dimensions with an index of refraction of approximately 1.5 were procured from Advanced Glass Industries. The Cherenkov light produced by electrically charged particles is detected by UV-glass window PMTs (model ET9814WB) and quartz window ET9814QB [ET_9814] photomultiplier tubes optically coupled to the quartz bars through RTV™615 silicon rubber of 50 m thickness. Thicker optical couplings resulted in fewer photoelectrons, so excellent quality control of the maximum thickness was required. The 16 bars in use in the hodoscope quartz plane are staggered in the beam direction with a transverse overlap between adjacent bars of 0.5 cm. The quartz plane frame allows for more bars to be added.
4.2.2 Calibration of Hodoscope
@@ 4.3 Drift Chambers
4.3.1 Design and Construction
The SHMS horizontal drift chambers provide information to determine the trajectory of charged particles passing through the detector stack. The drift chamber package consists of two horizontal drift chambers (DC1, DC2) separated by a distance of 1.1 m and oriented in the detector stack such that the sense wires planes are perpendicular to the central ray. Each chamber consists of a stack of six wire planes providing information on the track position along a single dimension in the plane of the wires and perpendicular to the wire orientations to better than 250 m. The perpendicular distance of the track relative to the wire is determined from the time of the signal produced by the ionization electrons as they drift from their production point to the wire in an electric field of approximately 3700 V/cm.
@@ 4.3.1 Design and Construction
Each chamber consists of two identical half chambers separated by a fiberglass mid-plane which supports the amplifier discriminator cards required for the sense wire readout. To minimize the production costs, only two unique PCB types were designed: an X-plane with wires oriented horizontally, and a U-plane with wires oriented at +60∘ relative the X-plane. All other plane orientations are generated by rotations of these two basic board types. For instance, the boards are designed such that a rotation of 180∘ in-plane about an axis through the center of the board produces boards with wires of the same orientation, but shifted by 1/2 cell width, thus allowing the resolution of left/right ambiguities. Rotation of the X-plane and the U-plane such that the top becomes the bottom produces the X’ and U’ orientations. The V and V’ boards, with wire orientation of -60∘ relative to the X-plane, are produced by a rotation of the U and U’ boards of 180∘ into the page about a vertical axis though the center of the board. Each half chamber has three planes with the first half consisting of (U, U’, X) and the second half consisting of (X’, V’, V). The first chamber, DC1, is oriented in the SHMS frame such that the board ordering as seen by particle traversing the spectrometer is (U, U’, X, X’, V’, V), while for the second chamber, DC2, the ordering is reversed (V, V’, X’, X, U’, U) as is shown in Figure 28.
@@ 4.3.2 Calibration
In the approximation that the incident particle distribution is uniform, or has a constant slope, a single cell111A cell is one sense wire surrounded by field wires such that the sense wire is at the center and the field wires are at the corners. on average sees a flat distribution of events. The measured drift time distribution can therefore be mapped to drift distance by areal scaling. This distribution can be determined for the individual wire or averaged over an entire group (up to 16 wires per discriminator card). Associated with each drift time distribution is a time, , which corresponds to the time at which ionized particles come into contact with the wire. If this value is non-zero, this is the value by which all drift times must be shifted in order to assure that . All subsequent times in each spectrum are measured relative to this time.
@@ 4.4 Heavy-Gas Cherenkov Counter
4.4.1 Design and Construction
The SHMS Heavy-Gas Cherenkov detector (HGC) is a threshold-type Cherenkov detector, designed to separate charged and over most of the SHMS operating momentum range, 3–11 GeV/c. The radiator gas C4F10 at 1 atm, with an index of refraction of =1.00143 at standard temperature [btev], allows to produce abundant Cherenkov light above 3 GeV/c momentum, while remain below Cherenkov threshold until about 7 GeV/c. Optimal separation at higher momenta requires a reduction in the gas pressure, down to 0.3 atm at 11 GeV/c.
A schematic view of the detector is shown in Fig. 30. The SHMS focal plane is subtended by four 5560 cm 0.3 cm thick glass mirrors, which reflect the Cherenkov radiation to four Hamamatsu R1584 12.5 cm diameter photomultiplier tubes located above and below the particle envelope. The mirrors and gas are enclosed in a cylindrical aluminum tank of 164.9 cm inner diameter and 113.5 cm length, with entrance and exit windows of hydroformed 0.102 cm thickness 2024 T-4 aluminum alloy [ASM]. The vessel is sufficiently strong to be pumped to vacuum before introducing the radiator gas, avoiding the need to purge when filling. A unique aspect of the detector is the placement of the photomultipliers outside the gas envelope, viewing the enclosure through 1.00 cm thick Corning 7980 quartz windows. This allows the gas enclosure to be smaller in diameter than would otherwise be possible, as the full length of the PMT and base no longer need to be fully within the diameter of the vessel. It also makes the PMTs available for servicing without venting the gas.
The mirrors are inexpensive, having been produced by the slumping process [sinclair]. As a result, they deviate from the desired 110 cm radius of curvature with a slightly oblate shape [wli_msc]. However, the Cherenkov cone on the mirrors for 3-7 GeV/c in C4F10 is 7-10 cm in diameter, so optical quality mirrors are not required for this application. The UV wavelength characteristics of the respective optical components are relatively well matched. C4F10 has good transmittance down to 160 nm [btev]. The quartz viewing windows provide 88% transmission down to 200 nm, including the 10% loss due to surface reflection [corning], and the optical glass faced PMTs have 70% of their peak quantum efficiency at 200 nm (peak at 350 nm) [hamamatsu]. Accordingly, the mirror reflectivity was optimized for 90% at 270 nm, and 75% at 200 nm [eci].
@@ 4.4.2 Calibration
where is the distance from the focal plane to the HGC mirror plane. The coordinate axis for the HGC is the convention used in charged particle transport in dispersive magnetic systems. The -axis is the direction of increasing particle momentum, the -axis is the direction of particle travel through the spectrometer, and the -axis is deduced from . Additionally, timing cuts are applied to the HGC data, collected using the high resolution pulse time setting in the FADC250’s FPGA [JLABF250]. The pulse time reconstruction in firmware corresponds to the time at which a pulse reaches half of its maximum amplitude after passing a threshold of 5 mV. Lastly, a cut on particle velocity, , is applied.
An example of a completed calibration is shown in Figs. 31, 32. For this run, the HGC was filled with C4F10 at 1 atm, and the SHMS central momentum was 2.583 GeV/c with polarity set to detect positively-charged particles. Cherenkov radiation is produced by traversing the HGC with momentum GeV/c. This can occur only for , which corresponds roughly to the bottom half of the HGC. Subthreshold with , as well as and , may produce low-level light in the HGC via knock-on electron emission and scintillation in the radiator gas. The adjacent mirror cuts described above produce a clear SPE peak in Fig. 31.
@@
An inherent systematic uncertainty is present in the HGC calibration due to statistical errors in determining the location of the SPE peak in the various mirror quadrants. This uncertainty was quantified by recording the locations of the SPE across several runs, for the different adjacent mirror combinations for each PMT, as well as by varying the contribution of the higher PE tail extending underneath the SPE peak, as in Figs. 31, 32. The systematic uncertainty in the calibration is taken to be the root mean square of this set of values, giving %. It should be noted this uncertainty is somewhat larger than the statistical uncertainty of the SPE peak, which is typically 0.2 to 0.6%.
@@ 4.5.1 Design and Construction
Four spherical thin glass mirrors of radius 135 cm, square in shape with edge lengths of 43 cm, focus the Cherenkov light onto the PMTs. The glass blanks were manufactured by Rayotek Scientific [Rayotek] from borosilicate glass of 3 mm thickness by slumping over a polished steel mold and then cutting to dimensions. Simulations showed a reduction of collection efficiency due to incoming photon losses at the exposed edges of the mirror. As such, the edges were bevelled away from the active surface to minimize scattering from these edges.
@@
The blanks were coated by the Thin Film and Glass Service of the Detector Technologies Group at CERN [DTG]. The reflectivity was also measured at CERN and found to be excellent well into the UV (Fig. 35).
Like the HGC, the four mirrors are arranged in a 2 by 2 array with a small overlap in the center, providing full coverage over the active area. In order to accomplish this without mechanical interference, the mirrors were staggered along the tank -axis which is the direction of the incoming particles. The mirrors were mounted in a monolithic frame installed as single unit (see Figure 36).
The four PMTs are 14 stage 5” quartz window PMTs manufactured by Electron Tubes Enterprises [ETTubes], model 9823QKB04. The 9823QKB04 has a quantum efficiency above 5% at 150 nm and 30% at 350 nm. The tubes are surrounded by a mu-metal shield and the positive HV is distributed to the stages by a resistive divider. After evidence of corona or arcing inside one of the bases due to the relatively low dielectric breakdown strength of the gas mixture, the problem was resolved by applying silicon conformal coating to suspect areas.
@@ 4.6 Aerogel Cherenkov Counter
4.6.1 Design and Construction
Fig. 38 shows a drawing of the aerogel counter installed downstream of the cylindrical HGC in the SHMS detector stack. The detector consists of two main components: a tray which holds the aerogel material, and a light diffusion box with photomultiplier tubes (PMTs) for light readout. Four identical trays for aerogel of nominal refractive indices of 1.030, 1.020, 1.015 and 1.011 were constructed. These allow for particle identification over the wide range of momenta summarized in Table 4. The design allows for easy detector assembly and replacement of the aerogel trays. The aerogel tray is 9,cm thickness and the total thickness of the detector is 24.5 cm along the optical axis of the SHMS. A detailed discussion of the detector, characterization of its components, and performance tests can be found in Refs. [AeroNIM17].
@@ 4.6.1 Design and Construction
The diffusion box is made of the aluminum alloy 6061-T6. The side panels are constructed of 2.5 cm (1-inch) plates. The back cover is 1 mm (1/16”) thick. The inner dimensions of the box are cm3 (40.5” 44.5” 6.82”). To optimize light collection, the inner surface of the diffusion box is lined with either 3 mm (covering 60% of the surface) or 1 mm (remaining 40% of the surface) thick GORE diffuse reflector material [Gore]. This material has a reflectivity of about 99% over the entire spectrum.
The light collection is handled by 5” diameter photomultiplier tubes (XP4500). The 5.56” (14.1 cm) diameter cylindrical housings holding the PMTs are mounted upon 14 waterjet cut circular openings on the left and right (long) sides of the diffusion box, with minimum spacing of 14.92 cm (5.875”) between the centers. The PMTs are sealed into their housing using a light-tight synthetic rubber material (Momentive RTV103 Black Silicone Sealant) and the whole assembly is sealed light-tight.
The trays contain two different types of aerogel. The aerogel tiles of refractive indices =1.030 and =1.020 were manufactured by Matsushita Electric Works, Ltd prior to 2010. The tiles of refractive indices =1.015 and =1.011 were manufactured by the Japan Fine Ceramics Center between 2010 and 2013. All tiles have dimensions of approximately 11 cm by 11 cm by 1.1 cm and are hydrophobic. The depth of the aerogel radiator in their trays is on average 9 cm thick (8 layers). For the SP-30, SP-20 and SP-15 trays the aerogel covers and area of 110 cm x 100 cm. In the SP-11 aerogel tray the radiator covers an active area of 90 cm x 60 cm. To improve the reflectivity, and thus light collection, inside the trays for the lowest two refractive indices 1 mm thick GORE diffusive reflector material (DRP-1.0-12x30-PSA) with reflectivity of about 99% was used. The SP-30 and SP-20 aerogel trays were covered with 0.45 m thick Millipore paper membrane GSWP-0010 (Millipore).
@@ 4.7 Preshower and Shower Counters
A nearly complete description of the design, pre-assembly component checkout, and construction of the SHMS calorimeter can be found in Ref. [mkrtchyan]. This section provides a brief summary.
4.7.1 Design and Construction